OpenStudy (atomicreaper):

If (2 − 3i) + (x + yi) = 6, what is x + yi?

OpenStudy (error1603):

k

OpenStudy (xmissalycatx):

Stop @Error1603

OpenStudy (xmissalycatx):

Do you have answer choices? XD

OpenStudy (jchick):

(2−3i)+(x+yi)=6 (2+x)+(−3+y)i=6+0i

OpenStudy (jchick):

Beginning with the given (2−3i)+(x+yi)=6 eliminate the parentheses. Rearrange the terms so that you have only x+yi on the left side.

OpenStudy (atomicreaper):

How?

OpenStudy (jchick):

Which part would you like me to explain more how what?

OpenStudy (jchick):

Oh how to remove the parenthesis?

OpenStudy (atomicreaper):

yes

OpenStudy (jchick):

Simplifying (2 + -3i) + (x + yi) = 6

OpenStudy (jchick):

Remove parenthesis around (2 + -3i) 2 + -3i + (x + yi) = 6

OpenStudy (jchick):

Reorder the terms: 2 + -3i + (iy + x) = 6

OpenStudy (jchick):

Just take them off

OpenStudy (jchick):

Remove parenthesis around (iy + x) 2 + -3i + iy + x = 6

OpenStudy (jchick):

Solving 2 + -3i + iy + x = 6

OpenStudy (atomicreaper):

Then what?

OpenStudy (jchick):

So now we will solve for i

OpenStudy (jchick):

Solving for variable 'i'. Move all terms containing i to the left, all other terms to the right. that is ok

OpenStudy (jchick):

Do you understand?

OpenStudy (jchick):

Add '-2' to each side of the equation. 2 + -3i + iy + -2 + x = 6 + -2

OpenStudy (jchick):

Reorder the terms: 2 + -2 + -3i + iy + x = 6 + -2

OpenStudy (atomicreaper):

ok then what?

OpenStudy (jchick):

Combine like terms: 2 + -2 = 0 0 + -3i + iy + x = 6 + -2 -3i + iy + x = 6 + -2 Hang on I am doing it as fast as I can

OpenStudy (jchick):

Combine like terms: 6 + -2 = 4 -3i + iy + x = 4

OpenStudy (jchick):

Add '-1x' to each side of the equation. -3i + iy + x + -1x = 4 + -1x

OpenStudy (jchick):

Combine like terms: x + -1x = 0 -3i + iy + 0 = 4 + -1x -3i + iy = 4 + -1x

OpenStudy (jchick):

Reorder the terms: -4 + -3i + iy + x = 4 + -1x + -4 + x Reorder the terms: -4 + -3i + iy + x = 4 + -4 + -1x + x

OpenStudy (jchick):

Combine like terms: 4 + -4 = 0 -4 + -3i + iy + x = 0 + -1x + x -4 + -3i + iy + x = -1x + x

OpenStudy (jchick):

Combine like terms: -1x + x = 0 -4 + -3i + iy + x = 0

OpenStudy (atomicreaper):

ok then what?

OpenStudy (jchick):

Well can you figure it out from here?

OpenStudy (atomicreaper):

yeah I got it thanks.

OpenStudy (jchick):

Ok awesome